Solve by the elimination method.
7r - 5s =-18
5r + 7s = 40
Help me please. This is getting sooooo much harder and I am grasping none of it.This is terrible!!!

Answer by scott8148(6628)

multiplying 1st equation by 7 __ 49r-35s=-126 __ because coefficient of s in 2nd equation is 7

multiplying 2nd equation by 5 __ 25r+35s=200 __ because cofficient of s in 1st equation is -5

adding the two new equations ELIMINATES the s terms __ 49r-35s+25r+35s=-126+200

74r=74 __ r=1

substituting __ 25(1)+35s=200 __ 35s=175 __ s=5